OpenStudy (anonymous):

4th root radical sign 2/3x^2

TheSmartOne (thesmartone):

\[\sqrt[4]{\frac{ 2 }{ 3 }x^2}\]

TheSmartOne (thesmartone):

Is that your question?

OpenStudy (anonymous):

its 3x^2 on the bottom

OpenStudy (jdoe0001):

hmm I don't think he ever asked anything, just a posting of a radical

TheSmartOne (thesmartone):

So is it \[\sqrt[4]{\frac{ 2 }{ 3x^2}}\]

OpenStudy (anonymous):

yes correct